OpenStudy (anonymous):

Solve for x in the proportion 9 over 27 times x equals x over the quantity x plus 3. Answer x ˜ -0.85 and x ˜ 1.18 x ˜ -6 and x ˜ 6.16 x ˜ -2.48 and x ˜ 2.48 x = 2 and x = 0

OpenStudy (anonymous):

9/27x = x/(x+3) 9(x+3) = 27x^2 0 = 27x^2 -9x - 27 0 = x^2 - (1/3)x - 1 1 = (x - 1/6)^2 - 1/36 37/36 = (x - 1/6)^2 sqrt(37/36) plus/minus 1/6 = x

OpenStudy (anonymous):

sorry plus minus sqrt(37/36) + 1/6 = x so x = -.85 or 1.18

OpenStudy (radar):

(9/27x)=x/(x+3) 27^2-9x-27=0 divide by 9 3x^2-x-3=0 Using quadratic equation: \[1\pm \sqrt{1 + 36}\over6\] First one is the solution